Solve 6(n+1 power)≥ 10. Round to four decimal places. This is Greek to me. I'm trying to help my niece and I have no clue where to start. Your assistance will be greatly appreciated.
Answer by scott8148(6628)

taking the log gives (n+1)(log6)>=(log10) ... n+1>=log10/log6 ... n>=(log10/log6)-1 ... n>=0.2851